### v1

@@ Abstract
Multicenter Lymphomas Benchmark, Multiple Instance Learning, Whole Slide Images, Pathology Foundation Models.
@@ 1 Introduction
Figure 1: A: Our multicenter lymphoma benchmark datasets comprises 5 classes : Follicular Lymphoma (FL), Chronic Lymphocytic Leukemia (CLL), Mantle Cell Lymphoma (MCL), Diffuse Large B-Cell Lymphoma (DLBCL) and healthy controls (NEG), with 999999 HE WSIs from 609609 individuals. B: Thumbnails of exemplary CLL WSIs. C: Random patches of $256\times 256$ pixels at $40\times$ magnification.
@@ 2.4 Lymphoma subtype classification
We use the deep multiple instance learning (MIL) framework for the trainable aggregation layer to classify the bag $\mathbf{H}$ of instance features into the five classes of our lymphoma dataset. The aggregation layer takes as input a bag $\mathbf{H}$ of $N$ instance features and outputs a patient latent vector $\mathbf{z}=\sum_{k=1}^{N}a_{k}\mathbf{h}_{k}$, where $a_{k}$ is the attention score. Finally, the patient latent vector $\mathbf{z}$ is passed to classifier to obtain predicted class scores $\mathbf{\hat{Y}}$. During training, the cross-entropy loss is computed between the predicted WSI class label $\mathbf{\hat{Y}}$ and ground-truth label $\mathbf{Y}\in\left\{1,\ldots,C\right\}$.
@@ 3.2 Evaluation
Magnification Encoder
Embedding
Dim
Aggregator AUC F1-score BACC
ResNet50
1024 AB-MIL 0.89 $\pm$0.03 0.66 $\pm$0.06 0.67 $\pm$0.06
@@
Titan
(Slide-level)
768 Linear 0.95 $\pm$0.01 0.78 $\pm$0.02 0.78 $\pm$0.01
@@
Titan
(Slide-level)
768 Linear 0.95 $\pm$0.01 0.77 $\pm$0.02 0.77 $\pm$0.02
@@
Titan
(Slide-level)
768 Linear 0.94 $\pm$0.01 0.75 $\pm$0.04 0.75 $\pm$0.04
@@ 3.3 Implementation details
The MIL models are trained under a weakly supervised learning setting, as only slide-level labels are available and no detailed annotations for individual patches. We use the MIL architecture, employing three variants: attention-based deep MIL (AB-MIL) [ilse2018attention], transformer-based MIL (TransMIL) [shao2021transmil], and TransMIL + BEL [sens2023bel]. These models aggregate features extracted from individual instances (patches) within each WSI to generate a single, comprehensive representation of the entire slide. By learning to assign different weights to instances, the architectures can capture the most relevant morphological features contributing to the classification task.
@@
Dataset Encoder
Embedding
Dim
Aggregator AUC F1-score BACC
@@
Encoder
OOD
Dataset
Aggregator AUC F1-score BACC
@@
Figure 4: Confusion matrix comparison on in-distribution Munich and out-of-distribution testsets: IID testset, Munich, and external OOD test: Kiel, Augsburg, and Erlangen. The MIL models perform strongest on Munich, and better results on Kiel and Erlangen, while lower perfomance on Augsburg due to noisy IHC stained labels.
@@ 3.4 Quantitative results
For lymphoma subtype prediction, we assessed model performance across varying magnifications, multiple patch-level and slide-level feature encoders, and both in-distribution and out-of-distribution (OOD) test sets.
Magnification study (Table 1): On the in-distribution Munich test set, we compared baseline ResNet50, the patch-level encoder UNI2, and the slide-level encoder Titan across 40×, 20×, and 10× magnifications. TransMIL combined with UNI2 or Titan is better than AB-MIL at 40× and 20×, with a slight performance decline observed at 10×. Despite this, AB-MIL remained the most computationally efficient option for training and inference compared to transformer-based MIL methods.
@@
Out-of-distribution generalization (Table 3): Models trained on the Munich dataset using UNI2 or Titan features were evaluated on three external cohorts: Kiel, Augsburg, and Erlangen. UNI2 paired with AB-MIL achieved the best overall performance on OOD Kiel and Augsburg in the multi-class setting. For the Erlangen dataset, which contains only CLL cases, UNI2 combined with TransMIL yielded the highest accuracy. Across all OOD evaluations, UNI2 + AB-MIL demonstrated the most robust and stable generalization behavior.
Confusion matrix comparison on in-distribution and out-of-distribution testsets: Figure 4 presents confusion matrices for UNI2 + AB-MIL on the in-distribution Munich test set and the three OOD cohorts. The model performed strongest on the Munich dataset, followed by solid results on Kiel and Erlangen. Performance on Augsburg was comparatively lower, particularly for CLL and FL, which we attribute to noisy labels—Augsburg subtype labels were derived from IHC staining without pathologist confirmation, leading to potential misclassifications.
@@ 3.5 Whole slide image attention visualization
The attention scores can be visualized as heatmaps to highlight diagnostically informative regions, areas assigned high attention, while de-emphasizing regions of low relevance, such as normal tissue or background artifacts. To interpret the spatial distribution of model attention across a WSI, we convert the attention scores corresponding to the predicted class into percentiles and map these normalized values back to their original coordinates on the slide. Fine-grained heatmaps are produced by extracting overlapping patches and averaging the attention values within the overlapping areas. In our workflow, we employ Trident [zhang2025standardizing] to generate WSI-level attention visualizations (Figure 5 and Figure 6) and to display the patches with the highest attention scores. We note that a high-attention score in CLL is given to non-lymphoma areas and in DLBCL to poorly preserved areas (Figure 5). This behavior might be avoidable when training the algorithm with more classes.
@@ 4 Limitations
Although this multicenter benchmark is comprehensive, several limitations remain. First, the substantial decrease in performance ($20\%$) on external test sets highlights unresolved out-of-distribution generalization challenges. Scanner-specific color variation, staining differences, and shifts in patient cohorts across clinical sites continue to affect model robustness, as reflected by the performance drop on OOD cohorts. Second, the Augsburg dataset contains subtype labels derived automatically from IHC staining without pathologist confirmation, introducing label noise that likely contributes to the lower classification accuracy observed in this cohort. Finally, some cohorts exhibit class imbalance, particularly for rare subtypes or limited-control classes, which may bias model training and lead to overconfident predictions for majority classes.
@@ 5 Conclusion
On in-distribution test sets, all five foundation models (H-optimus-1, H0-mini, Virchow2, UNI2, Titan) achieved comparable performance with balanced accuracies exceeding $80\%$, demonstrating that current pathology foundation models have reached a similar level of representation quality for lymphoma subtyping. Both attention-based (AB-MIL) and transformer-based (TransMIL) aggregators performed similarly, with AB-MIL being computationally more efficient. Our magnification study revealed that 40× resolution is sufficient for accurate classification, with no additional benefit from higher resolutions or cross-magnification aggregation strategies. Despite strong in-distribution performance, we observed substantial performance drops on out-of-distribution test sets (balanced accuracy around $60\%$), revealing significant generalization challenges. This degradation likely stems from scanner-specific color variations, site-specific staining protocols, and differences in patient populations across centers. The Augsburg cohort, where labels were derived automatically from IHC staining without pathologist confirmation, showed particularly poor performance, highlighting the critical importance of high-quality ground truth annotations for model development. Additionally, we did not assess computational requirements or inference times, which are crucial considerations for clinical deployment. In contrast to expensive and time-consuming IHC staining, flow cytometry, and molecular genetic tests, our approach demonstrates that inexpensive HE-stained WSIs can accurately determine common lymphoma subtypes when evaluated within the same clinical environment. An AI-augmented pathology workflow could guide initial screening to prioritize cases requiring expert review, potentially reducing turnaround times and allowing more efficient use of specialized diagnostic tests.

### v2

@@ Abstract
Multicenter Lymphoma Benchmark, Multiple Instance Learning, Whole Slide Images, Pathology Foundation Models.
editors: Under Review for MIDL 2025
@@ 1 Introduction
Figure 1: A: Our multicenter lymphoma benchmark datasets comprises 5 classes : Follicular Lymphoma (FL), Chronic Lymphocytic Leukemia (CLL), Mantle Cell Lymphoma (MCL), Diffuse Large B-Cell Lymphoma (DLBCL) and healthy controls (NEG), with total 999999 HE WSIs from $609$ individual patients. The total number of WSIs in each class are given. B: Thumbnails of exemplary CLL WSIs. C: Random patches of $256\times 256$ pixels at $40\times$ magnification.
@@ 2.4 Lymphoma subtype classification
We use the deep multiple instance learning (MIL) framework for the trainable aggregation layer to classify the bag $\mathbf{H}$ of instance features into the five classes of our lymphoma dataset. The aggregation layer takes as input a bag $\mathbf{H}$ of $N$ instance features and outputs a patient latent vector $\mathbf{z}=\sum_{k=1}^{N}a_{k}\mathbf{h}_{k}$, where $a_{k}$ is the attention score.
Finally, the patient latent vector $\mathbf{z}$ is passed to classifier to obtain predicted class scores $\mathbf{\hat{Y}}$. During training, the cross-entropy loss is computed between the predicted WSI class label $\mathbf{\hat{Y}}$ and ground-truth label $\mathbf{Y}\in\left\{1,\ldots,C\right\}$.
@@ 3.2 Evaluation
Magnification Encoder Embedding Dim Aggregator AUC F1-score BACC
ResNet50 (Beseline) 1024 AB-MIL 0.89 $\pm$0.03 0.66 $\pm$0.06 0.67 $\pm$0.06
@@
Titan (Slide-level) 768 Linear 0.95 $\pm$0.01 0.78 $\pm$0.02 0.78 $\pm$0.01
@@
Titan (Slide-level) 768 Linear 0.95 $\pm$0.01 0.77 $\pm$0.02 0.77 $\pm$0.02
@@
Titan (Slide-level) 768 Linear 0.94 $\pm$0.01 0.75 $\pm$0.04 0.75 $\pm$0.04
@@ 3.3 Implementation details
The MIL models are trained under a weakly supervised learning setting, as only slide-level labels are available and no detailed annotations for individual patches. We use the MIL architecture, employing three variants: attention-based deep MIL (AB-MIL) [ilse2018attention], transformer-based MIL (TransMIL) [shao2021transmil], and TransMIL + BEL [sens2023bel]. These models aggregate features extracted from individual instances (patches) within each WSI to generate a single, comprehensive representation of the entire slide. By learning to assign different weights to instances, the architectures can capture the most relevant morphological features contributing to the classification task. We address class imbalance by applying balanced sampling at the patient level during aggregator model training, rather than relying on slide-level label balancing. This strategy mitigates subtype imbalance effects and ensures that model performance is not biased by overrepresented classes.
@@
Dataset Encoder Embedding Dim Aggregator AUC F1-score BACC
@@
Encoder OOD Dataset Aggregator AUC F1-score BACC
@@
Figure 4: Confusion matrix comparison on in-distribution Munich and out-of-distribution testsets: IID testset, Munich, and external OOD test: Kiel, Augsburg, and Erlangen. The MIL models perform strongest on Munich, and better results on Kiel, Augsburg, and Erlangen.
@@ 3.4 Quantitative results
For lymphoma subtype prediction, we assessed model performance across varying magnifications, multiple patch-level and slide-level feature encoders, and both in-distribution and out-of-distribution (OOD) test sets. Our lymphoma subtyping findings should be interpreted as complementary to the “all-task” benchmark [bareja2025evaluating] rather than directly comparable.
Magnification study (Table 1): On the in-distribution Munich test set, we compared baseline ResNet50, the patch-level encoder UNI2, and the slide-level encoder Titan across 40×, 20×, and 10× magnifications. TransMIL combined with UNI2 or Titan is better than AB-MIL at 40× and 20×, with a slight performance decline observed at 10×. Despite this, AB-MIL remained the most computationally efficient option for training and inference compared to transformer-based MIL methods. TransMIL + BEL show only minimal gains over TransMIL. This behavior is likely due to a mismatch between the BEL objective and the morphological characteristics of lymphoma WSIs: While BEL is designed to regularize bag-level embeddings in transformer-based MIL, lymphoma subtyping relies on heterogeneous and spatially dispersed patterns rather than a single dominant bag embedding. We believe that TransMIL already captures the discriminative context in the slides, leaving limited room for additional gains from BEL. We trained TransMIL + BEL using the original hyperparameters proposed by [sens2023bel], but checked that further parameter tuning did not lead to improvements.
@@
Out-of-distribution generalization (Table 3): Models trained on the Munich dataset using UNI2 or Titan features were evaluated on three external cohorts: Kiel, Augsburg, and Erlangen.
UNI2 paired with AB-MIL achieved the best overall performance on OOD Kiel. The superior performance of UNI2 + AB-MIL on the OOD Kiel cohort is primarily attributable to UNI2’s stronger patch-level representations for cross-domain generalization. UNI2 is trained on large-scale, diverse histopathology datasets and preserves fine-grained cellular morphology, which is critical for lymphoma subtyping. In contrast, Titan operates at the slide level by aggressively reducing token resolution, which improves efficiency but limits its ability to capture subtle, spatially localized lymphoma patterns under domain shift. As a result, UNI2 features generalize more robustly to unseen centers.
UNI2 paired with AB-MIL achieved the best overall performance on OOD Augsburg. We use Augsburg data to assess model robustness under domain shift. To address the label-space mismatch, we report four-class metrics using the original predictions, considering the top-2 class for the those cases where NEG is the top-1 prediction, resulting in comparable performance of our models on the Augsburg dataset.
For the Erlangen dataset, which contains only CLL cases, UNI2 combined with TransMIL yielded the highest accuracy. In the Erlangen cohort, the high standard deviation observed is primarily due to its very small sample size (29 WSIs) and its single-class composition (CLL only). In this setting, balanced accuracy becomes highly sensitive to fold-level variations, leading to inflated variance. Additionally, Erlangen data were acquired using a different scanner, further amplifying instability. We report these results transparently to highlight the challenges of evaluating OOD performance on small, single-class cohorts, rather than as evidence of reliable generalization. Across all OOD evaluations, UNI2 + AB-MIL demonstrated the most robust and stable generalization behavior.
Confusion matrix comparison on in-distribution and out-of-distribution testsets: Figure 4 presents confusion matrices for UNI2 + AB-MIL on the in-distribution Munich test set and the three OOD cohorts. The model performed strongest on the IID Munich dataset, followed by solid results on Kiel, Augsburg, and Erlangen.
Computational efficiency: Table 4 compares computational efficiency of the five pathology foundation models, reporting parameters, FLOPs, and throughput (images per second) measured on a consumer-grade RTX 3090 GPU. For the patch-level FMs, throughput is computed in FP16 precision with a batch size of 32, and we report the mean and standard deviation over 500 batches. In whole slide image analysis, where thousands of image patches are encoded for a slide level prediction, this efficiency translates into total processing time with varying number of image patches per slide. For slide-level FMs such as Titan, it operates at the whole slide level by substantially reducing the number of processed tokens during inference. The throughput is computed as slides per second with 10,000 patches per slide on A100 GPU.
Table 4: Computational efficiency statistics of five pathology foundation models, including parameter count, FLOPs, and inference time per image or per slide. Throughput is computed on images per second for patch-level models, and slides per second for slide-level model (Titan).
Model Params. FLOPs Throughput
[M] $(\downarrow)$ [G] $(\downarrow)$ [img or slide/s] $(\uparrow)$
H-optimus-1 1,135 591.81 82.5 $\pm$4.3
H0-mini 87 44.60 885.1 $\pm$7.1
Virchow2 632 329.11 134.3 $\pm$7.9
UNI2 681 360.71 134.6 $\pm$7.6
Titan 85 36.76 43.95 $\pm$2.3
@@ 3.5 Whole slide image attention visualization
The attention scores can be visualized as heatmaps to highlight diagnostically informative regions, areas assigned high attention, while de-emphasizing regions of low relevance, such as normal tissue or background artifacts. To interpret the spatial distribution of model attention across a WSI, we convert the attention scores corresponding to the predicted class into percentiles and map these normalized values back to their original coordinates on the slide. Fine-grained heatmaps are produced by extracting overlapping patches and averaging the attention values within the overlapping areas. In our workflow, we employ Trident [zhang2025standardizing] to generate WSI-level attention visualizations (Figure 5 and Figure 7) and to display the patches with the highest attention scores. In Figure 6, UNI2 features generalize more robustly than Titan, and avoid artifact patches. We note that a high-attention score in CLL is given to non-lymphoma areas and in DLBCL to poorly preserved areas (Figure 5). This behavior might be avoidable when training the algorithm with more classes.
Figure 6: Attention visualization of CLL lymphoma subtype on OOD Erlangen cohort. UNI2 avoids artifact patches in the highest attention scores as compared to Titan.
@@ 4 Limitations
Although this multicenter benchmark is comprehensive, several limitations remain. First, the substantial decrease in performance ($20\%$) on external test sets highlights unresolved out-of-distribution generalization challenges. Scanner-specific color variation, staining differences, and shifts in patient cohorts across clinical sites continue to affect model robustness, as reflected by the performance drop on OOD cohorts. Second, some cohorts exhibit class imbalance, particularly for rare subtypes or limited-control classes, which may bias model training and lead to overconfident predictions for majority classes.
@@ 5 Conclusion
Current “all-task” benchmark do not include lymphoma subtyping, and TCGA offers only a single lymphoma subtype (DLBCL) with limited histological diversity. In contrast, our multicenter benchmark focuses specifically on lymphoma, covering four common subtypes and healthy controls to enable a focused evaluation of this clinically important but so far underrepresented disease. On in-distribution test sets, all five foundation models (H-optimus-1, H0-mini, Virchow2, UNI2, Titan) achieved comparable performance with balanced accuracies exceeding $80\%$, demonstrating that current pathology foundation models have reached a similar level of representation quality for lymphoma subtyping. Both attention-based (AB-MIL) and transformer-based (TransMIL) aggregators performed similarly, with AB-MIL being computationally more efficient. Our magnification study revealed that 40× resolution is sufficient for accurate classification, with no additional benefit from higher resolutions or cross-magnification aggregation strategies. Despite strong in-distribution performance, we observed substantial performance drops on out-of-distribution test sets (balanced accuracy around $60\%$), revealing significant generalization challenges. This degradation likely stems from scanner-specific color variations, site-specific staining protocols, and differences in patient populations across centers. In contrast to expensive and time-consuming IHC staining, flow cytometry, and molecular genetic tests, our approach demonstrates that inexpensive HE-stained WSIs can accurately determine common lymphoma subtypes when evaluated within the same clinical environment. An AI-augmented pathology workflow could guide initial screening to prioritize cases requiring expert review, potentially reducing turnaround times and allowing more efficient use of specialized diagnostic tests.